# A single-loop SPIDER-type stochastic subgradient method for expectation-constrained nonconvex nonsmooth optimization

### v1

@@ Appendix A Comparisons among our CQs and others
(50) $\lambda^{+}\left(g(\mathbf{x})-g\left(\mathbf{x}^{+}\right)\right)\geq\lambda^{+}\left(\left\langle\boldsymbol{\zeta}_{g},\mathbf{x}-\mathbf{x}^{+}\right\rangle-\frac{\rho_{g}}{2}\|\mathbf{x}-\mathbf{x}^{+}\|^{2}\right)$
$=$ $\left\langle\lambda^{+}\boldsymbol{\zeta}_{g},\mathbf{x}-\mathbf{x}^{+}\right\rangle-\frac{\lambda^{+}\rho_{g}}{2}\|\mathbf{x}-\mathbf{x}^{+}\|^{2}=\left\|\mathbf{x}-\mathbf{x}^{+}\right\|^{2}-\frac{\lambda^{+}\rho_{g}}{2}\|\mathbf{x}-\mathbf{x}^{+}\|^{2}.$
Since $0<g({\mathbf{x}})\leq l_{g}\textnormal{dist}\left(\mathbf{x},\Omega\right)\leq B_{g}$, dividing both sides of Eq. 50 by $\lambda^{+}$ yields
@@
$g(\mathbf{x})-g\left(\mathbf{x}^{+}\right)$ $\geq\frac{\left\|\mathbf{x}-\mathbf{x}^{+}\right\|^{2}}{\lambda^{+}}-\frac{\rho_{g}}{2}\|\mathbf{x}-\mathbf{x}^{+}\|^{2}=\|\boldsymbol{\zeta}_{g}\|\|\mathbf{x}-\mathbf{x}^{+}\|-\frac{\rho_{g}}{2}\|\mathbf{x}-\mathbf{x}^{+}\|^{2}$
$\geq\sqrt{2B(\overline{\rho}-\rho_{g})}\left\|\mathbf{x}-\mathbf{x}^{+}\right\|-\frac{\rho_{g}}{2}\|\mathbf{x}-\mathbf{x}^{+}\|^{2},$
where the last inequality follows from Lemma A.1. Noting $g(\mathbf{x})=g_{+}(\mathbf{x})$ for $\mathbf{x}\not\in\Omega$, $g\left(\mathbf{x}^{+}\right)=0$ and $\|\mathbf{x}-\mathbf{x}^{+}\|=\textnormal{dist}({\mathbf{x}},\Omega)\leq\frac{\sqrt{2B(\overline{\rho}-\rho_{g})}}{\rho_{g}}$, we obtain $\gamma\textnormal{dist}({\mathbf{x}},\Omega)\leq g_{+}({\mathbf{x}})$ and complete the proof.

### v2

@@ Acknowledgement
$\quad+\sum_{i\in\mathcal{J}_{2}}\lambda_{i}\left(\textstyle g_{i}(\mathbf{x})+\frac{\overline{\rho}}{2}\|\mathbf{x}-\mathbf{x}\|^{2}\right)+\left\langle\textstyle\sum_{i\in{\mathcal{J}}_{2}}\lambda_{i}{\boldsymbol{\zeta}}_{g_{i}},\mathbf{y}-\mathbf{x}\right\rangle$
(43) $\leq$ $\sum_{i\in\mathcal{J}_{1}\cap{\mathcal{J}}_{3}}\left(g_{i}(\mathbf{y})+\frac{\overline{\rho}}{2}\|\mathbf{y}-\mathbf{x}\|^{2}\right)+\sum_{i\in\mathcal{J}_{2}}\lambda_{i}\left(g_{i}(\mathbf{y})+\frac{\overline{\rho}}{2}\|\mathbf{y}-\mathbf{x}\|^{2}\right)\leq-B,$
where the first equality follows from $g_{i}(\mathbf{x})]_{+}=g_{i}(\mathbf{x}),\forall\,i\in\mathcal{J}_{1}\cap{\mathcal{J}}_{3},$ and $g_{i}(\mathbf{x})]_{+}=0$, ${\boldsymbol{\zeta}}_{g_{i}}=\mathbf{0},\forall\,i\in\mathcal{J}_{1}\cap\mathcal{J}_{3}^{c}$, the first inequality comes from the $(\overline{\rho}-\rho_{g})$-strongly convexity of $\sum_{i\in\mathcal{J}_{1}\cap{\mathcal{J}}_{3}}(g_{i}(\cdot)+\frac{\overline{\rho}}{2}\|\cdot-\mathbf{x}\|^{2})+\iota_{\mathcal{X}}(\cdot)$ and the convexity of $\sum_{i\in\mathcal{J}_{2}}\lambda_{i}\big{(}g_{i}(\cdot)+\frac{\overline{\rho}}{2}\|\cdot-\mathbf{x}\|^{2}\big{)}$, as ${\mathcal{J}}_{1}\cap{\mathcal{J}}_{3}\neq\emptyset$ and $\lambda_{i}\geq 0,\forall\,{i\in J_{2}}$, the last inequality holds by ${\mathcal{J}}_{1}\cap{\mathcal{J}}_{3}\neq\emptyset$, $\lambda_{i}\geq 0,\forall\,{i\in J_{2}}$, and $g_{i}(\mathbf{y})+\frac{\overline{\rho}}{2}\|\mathbf{y}-\mathbf{x}\|^{2}\leq-B$ for all $i\in[m]$.
The inequality A.1, together with Young’s inequality, gives
@@
$-\frac{\left\|\sum_{i\in J_{1}}{\boldsymbol{\zeta}}_{g_{i}}+{\mathbf{u}}+\sum_{i\in{\mathcal{J}}_{2}}\lambda_{i}\boldsymbol{{\boldsymbol{\zeta}}}_{g_{i}}\right\|^{2}}{2(\overline{\rho}-\rho_{g})}$
$\textstyle\leq$ $\left\langle\sum_{i\in J_{1}}{\boldsymbol{\zeta}}_{g_{i}}+\sum_{i\in{\mathcal{J}}_{2}}\lambda_{i}\partial g_{i}({\mathbf{x}})+{\mathbf{u}},\mathbf{y}-\mathbf{x}\right\rangle+\frac{\overline{\rho}-\rho_{g}}{2}\|\mathbf{y}-\mathbf{x}\|^{2}$
$\textstyle\leq$ $-B-\sum_{i\in\mathcal{J}_{1}}[g_{i}(\mathbf{x})]_{+}-\sum_{i\in\mathcal{J}_{2}}\lambda_{i}g_{i}(\mathbf{x})\leq-B,$
where the last inequality holds by $g_{i}({\mathbf{x}})\geq 0,\forall\,i\in\mathcal{J}_{1}\cap{\mathcal{J}}_{3}$, and $\lambda_{i}g_{i}({\mathbf{x}})\geq 0,\forall\,i\in J_{2}$. Since the choice of subgradients and $\mathbf{u}$} is arbitrary, the proof is completed.
Appendix B Problems 2–3 in Section 4.1 satisfy 2
We first verify in this section that the fairness-constrained logistic‐loss problem in Section 4.1 satisfies 2. Let $\rho_{g}$ and $l_{g}$ denote the weak convexity modulus and Lipschitz continuity modulus of $\mathbf{g}$} in Problem 2, respectively. Under mild conditions, the following lemma establishes the Slater-type CQ.
Lemma B.1 ([27, Theorem B.1]).
[…]
$q({\mathbf{x}}):=\max\left\{\max_{i=1,\dots,n_{\mathrm{p}}}\!\frac{\sigma({\mathbf{x}}^{\top}{\mathbf{a}}_{i}^{\mathrm{p}})(1-\sigma({\mathbf{x}}^{\top}{\mathbf{a}}_{i}^{\mathrm{p}}))}{n_{\mathrm{p}}},\max_{i=1,\dots,n_{\mathrm{u}}}\!\frac{\sigma({\mathbf{x}}^{\top}{\mathbf{a}}_{i}^{\mathrm{u}})(1-\sigma({\mathbf{x}}^{\top}{\mathbf{a}}_{i}^{\mathrm{u}}))}{n_{\mathrm{u}}}\right\}.$
Then the following statements hold:
  • $\mathrm{(i)}$
    Let $q_{\mathrm{min}}=\inf\{q(\mathbf{x})||\overline{\Psi}(\mathbf{x})\mid\in[(\kappa+\underline{\kappa})/2,(\kappa+\bar{\kappa})/2]\}.$ It holds $q_{\mathrm{min}}>0$.
  • $\mathrm{(ii)}$
    Let $\underline{q}\in(0,q_{\mathrm{min}}]$. Then Problem 2 in Section 4.1 satisfies the Slater-type CQ with any $\bar{\rho}>\rho_{g}$, and with parameters $0<B\leq\min\left\{\frac{\kappa-\underline{\kappa}}{2},\frac{\underline{q}^{2}}{\rho_{g}+4\bar{\rho}},\frac{\underline{q}^{2}\kappa}{l_{g}^{2}}\right\}$ and $0<B_{g}\leq\min\left\{\frac{\bar{\kappa}-\kappa}{2},\frac{\underline{q}^{2}}{\rho_{g}+4\bar{\rho}},\frac{\underline{q}^{2}\kappa}{l_{g}^{2}}\right\}.$
We note that though $q_{\mathrm{min}}$ is difficult to compute, for our tested Problem 2, we can directly establish a positive lower bound on $q({\mathbf{x}})$ uniformly about ${\mathbf{x}}\in{\mathcal{X}}$. In particular, let $D_{a}=\min\left(\min_{1\leq i\leq n_{\mathrm{p}}}\|{\mathbf{a}}_{i}^{\mathrm{p}}\|,\min_{1\leq j\leq n_{\mathrm{u}}}\|{\mathbf{a}}_{j}^{\mathrm{u}}\|\right)$, i.e., there is ${\mathbf{a}}_{i}^{\mathrm{p}}$ or ${\mathbf{a}}_{j}^{\mathrm{u}}$ such that $\|{\mathbf{a}}_{i}^{\mathrm{p}}\|=D_{a}$ or $\|{\mathbf{a}}_{j}^{\mathrm{u}}\|=D_{a}$. Since $\|{\mathbf{x}}\|\leq D,\forall\,{\mathbf{x}}\in{\mathcal{X}}$, it holds $|{\mathbf{x}}^{\top}{\mathbf{a}}^{\mathrm{p}}_{i}|\leq D_{a}D$ or $|{\mathbf{x}}^{\top}{\mathbf{a}}^{\mathrm{u}}_{j}|\leq D_{a}D.$ Consequently, we have $q({\mathbf{x}})\geq\frac{\sigma(D_{a}D)\bigl{(}1-\sigma(D_{a}D)\bigr{)}}{\max\{n_{\mathrm{p}},\,n_{\mathrm{u}}\}}=:\tilde{q},\forall\,{\mathbf{x}}\in{\mathcal{X}}$, and thus $q_{\mathrm{min}}\geq\tilde{q}$. Therefore, we can set $\underline{q}=\tilde{q}$ in the above Statement (ii).
Next, we verify that Problem 3 defined in Section 4.1 also satisfies 2. For each $i\in\{2,\ldots,M\}$, let $g_{i-1}$ be the component constraint function that corresponds to ${\mathcal{D}}_{i}$ in Problem 3, and suppose that ${\mathbf{a}}_{i}]_{1}=1,\forall\,{\mathbf{a}}_{i}\in{\mathcal{D}}_{i}$. Under this assumption, one can easily show that for any $i\in\{2,3,\ldots,M\}$,
(44) $\|\nabla g_{i-1}(\mathbf{x})\|$ $\geq\frac{1}{\left|\mathcal{D}_{i}\right|}\sum^{M}_{\begin{subarray}{c}l=2\\ l\neq i\end{subarray}}\sum_{{\mathbf{a}}_{i}\in\mathcal{D}_{i}}\sigma\left(\mathbf{x}_{i}^{\top}{\mathbf{a}}_{i}-\mathbf{x}_{l}^{\top}{\mathbf{a}}_{i}\right)\left(1-\sigma\left(\mathbf{x}_{i}^{\top}{\mathbf{a}}_{i}-\mathbf{x}_{l}^{\top}{\mathbf{a}}_{i}\right)\right)\left[\mathbf{a}_{i}\right]_{1},$
by the fact $\|\nabla g_{i-1}(\mathbf{x})\|\geq\|\nabla_{[{\mathbf{x}}_{i}]_{1}}g_{i-1}(\mathbf{x})\|$. Let $D_{a}=\max_{i=2,3,\ldots,M}\{\min_{{\mathbf{a}}_{i}\in{\mathcal{D}}_{i}}\|{\mathbf{a}}_{i}\|\}$, i.e., for each $i\in\{2,3,\ldots,M\}$, there is ${\mathbf{a}}_{i}\in{\mathcal{D}}_{i}$ such that $\|{\mathbf{a}}_{i}\|\leq D_{a}$. Since we have $\|{\mathbf{x}}\|\leq D,\forall\,{\mathbf{x}}\in{\mathcal{X}}$, it holds $|\mathbf{x}_{i}^{\top}{\mathbf{a}}_{i}-\mathbf{x}_{l}^{\top}{\mathbf{a}}_{i}|\leq 2D_{a}D$. By an analogous argument, we establish that for all $i\in\{2,3,\ldots,M\}$,
$\|\nabla g_{i-1}(\mathbf{x})\|\geq\frac{\sigma(2D_{a}D)\bigl{(}1-\sigma(2D_{a}D)\bigr{)}}{\max_{i=2,3,\ldots,M}\{|\mathcal{D}_{i}|\}}=:\tilde{q}$.
Hence, following steps (83)–(85) in the proof of [27, Theorem B.1], we are able to show that Problem 3 satisfies the Slater-type CQ as defined in 2. We do not repeat the details here.